Asanji and Ndiba each improved their yards by planting rose bushes and ornamental grass, They bought their supplies from the same store. Asanji spent $11 on 1 rose bush and 2 bunches of ornamental grass. Ndiba spent $23 on 5 rose bushes and 2 bunches of ornamental grass. Find the cost of one rose bush and the cost of one bunch of oranamental grass.

Answer by ikleyn(53763)

.
Asanji and Ndiba each improved their yards by planting rose bushes and ornamental grass,
They bought their supplies from the same store.
Asanji spent $11 on 1 rose bush and 2 bunches of ornamental grass.
Ndiba spent $23 on 5 rose bushes and 2 bunches of ornamental grass.
Find the cost of one rose bush and the cost of one bunch of oranamental grass.
~~~~~~~~~~~~~~~


This problem is for MENTAL solution.

The difference between their purchases is 4 rose bushes and 23-11 = 12 dollars.

Hence, one rose bush costs 12/4 = 3 dollars.

Then 2 bushes of ornamental grass costs 11 - 3 = 8 dollars --- hence, one
bush of ornamental grass costs 8/2 = 4 dollars.


Solved.